The population average fuel efficiency of U.S. light vehicles for 2005



was 21 mpg. If the standard deviation of the population was 2.9 and



the gas ratings were normally distributed, what is the probability that



the mean for a random sample of 25 light vehicles is under 20 mpg?



(Must show your work)



1
Expert's answer
2021-12-15T16:08:11-0500

Let "X=" the mean for a random sample: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=21mpg, \sigma=2.9mpg, n=25."


"P(X<20)=P(Z<\dfrac{20-\mu}{\sigma/\sqrt{n}})"

"=P(Z<\dfrac{20-21}{2.9/\sqrt{25}})\approx P(Z<-1.724138)"

"\approx 0.042341"